# Understanding and Characterizing Obfuscated Funds Transfers in Ethereum Smart Contracts

### v1

@@ 7. Impact on Existing Tools
Non-obfuscated 361 287 74 0.79 0.79 0.88
Obfuscated 92 11 81 0.12 0.11 0.21
Evaluation Results. Overall, its accuracy is 65.41% for all 453 samples. This contrasts with the higher precision reported in the original paper, suggesting that in real-world scenarios (especially for contracts with deeper obfuscation), many false negatives and positives exist. A closer look at the results, shown in Table 7, we can see that the effectiveness difference between obfuscated and non-obfuscated samples is very significant. SourceP achieves an F-1 score of 0.88 for non-obfuscated samples, which roughly aligns with the detection capability claimed in the original paper. However, for obfuscated contracts, both accuracy and recall drop sharply, with an F1 score of only around 0.21, indicating that obfuscation techniques in real-world Ponzi samples can remarkably affect the performance of state-of-the-art malware detection tools.
@@
Summary. Our study results demonstrate that obfuscation has a significant detrimental impact on existing detection tools (such as SourceP), causing a substantial drop in both accuracy and recall. In particular, for Ponzi schemes with obfuscation, the false negative rate can exceed 80%, rendering state-of-the-art detection tools useless. At a practical level, this observation underscores the importance of developing practical obfuscation analysis techniques, which is essential to mitigating the emerging security threats.
8. Discussion
8.1. Challenges and Future Work
Obfuscation techniques in smart contracts present significant challenges for auditing and regulatory practices, particularly in scam contracts, MEV bots, and highly centralized systems. These challenges include poor code readability, failure of conventional static detection tools, and delays in regulatory response. Several improvements are necessary to address these issues.
Enhanced Detection Techniques. Static analysis tools must be enhanced to track deeper control and data flows, while de-obfuscation preprocessing can simplify bytecode for more efficient audits. Dynamic analysis, such as runtime tracing or fuzzing, can bypass superficial obfuscation and validate fund flows during contract execution.
De-Obfuscation Strategies. Techniques from traditional software security, such as CFG flattening reversal and semantic normalization, can be adapted to EVM bytecode to identify critical logic like transfers and permission checks that obfuscation tries to hide.
Community-Driven Auditing. Establishing collaborative platforms, such as a ”contract blacklist” or ”high-risk obfuscation” repository, would allow researchers, auditors, and the public to tag suspicious contracts, improving transparency and enhancing collective oversight of the DeFi ecosystem.
8.2. Limitations
Our work has the following limitations. First, our findings are based on the Ethereum mainnet and do not cover other EVM-compatible blockchains like BSC, Polygon, or Tron. Hence, these blockchains may exhibit different obfuscation patterns. Second, our study only considers contracts deployed between 2021 and 2024. Thus, new compiler features or obfuscation techniques may emerge in future versions of Solidity or other blockchain platforms. Third, not all obfuscation strategies were covered. More advanced techniques, such as inline assembly abuse or internal EVM feature manipulation, require further investigation.
9. Related Work
@@
Smart Contract Security Analysis. Some research employs static analysis to enhance smart contract security and efficiency. USCHUNT (Bodell III et al., 2023) explores the balance between adaptability and security in upgradeable contracts. Madmax (Grech et al., 2018) targets vulnerabilities to prevent execution failures, while Slither (Slither, 2024) and Smartcheck (Tikhomirov et al., 2018) automatically detect flaws in Solidity contracts. Symbolic execution is also used to improve security; Mythril (ConsenSys, 2022) analyzes EVM bytecode, EthBMC (Frank et al., 2020) combines symbolic execution with concrete validation, and Reguard (Liu et al., 2018) and Manticore (of Bits, 2024) identify reentrancy and other bugs. Smartian (Choi et al., 2021) integrates fuzzing with static and dynamic analysis, while Confuzzius (Ferreira Torres et al., 2021) leverages data dependency insights for fuzzing. CRYSOL (Zhang et al., 2024) applies fuzzing to detect cryptographic defects in contracts. ContractFuzzer (Jiang et al., 2018) and Sfuzz (Nguyen et al., 2020) apply fuzzing to uncover security issues.
Research highlights various formal verification methods to enhance smart contract security. Sailfish (Bose et al., 2022) improves state inconsistency detection, while VetSC.(Duan et al., 2022) extends DApp verification. Zeus(Kalra et al., 2018) and Verx (Permenev et al., 2020) focus on contract safety and condition verification. Smartpulse (Stephens et al., 2021) analyzes time-based properties, Securify (Tsankov et al., 2018) identifies security breaches, and Verismart (So et al., 2020) ensures contract safety.
Excessive Owner Control. Centralization risk has been emphasized by Lamby et al. (2023), who reported that centralized backdoors were responsible for \$1.3 billion in DeFi losses in 2021 (Lamby et al., 2023). Lin et al. (2023) further identified centralization flaws, such as a single owner having the ability to modify transfer fees or update proxy contracts arbitrarily (Lin et al., 2024). Similarly, Yu et al. (2025) highlighted hidden backdoors, including arbitrary fund transfers (Yu and Lee, 2025). Another form of centralization, adjustable transaction fees, is frequently exploited by developers to impose excessive taxes on sales, as noted by Shiaeles and Li (2024) (Hall et al., 2024). Liquidity scams also leverage centralization, with advanced schemes using disguised function names or dynamic withdrawal logic to evade detection (Huynh et al., 2024; Lin et al., 2024).
Ponzi Schemes. Ponzi schemes on the blockchain gained prominence with the advent of Ethereum, prompting significant research into their detection. Galletta et al. (2024) compiled thousands of Ponzi scheme contract samples and developed classifiers to identify these schemes, focusing on investment-dividend models where profits rely on funds from later participants (Galletta and Pinelli, 2024). While detection methods for traditional Ponzi schemes are well-established, research on emerging DeFi variants remains limited. Earlier studies primarily addressed simple fund-sharing contracts, but newer schemes, such as Forsage, incorporate multi-level referral reward systems and cyclic dependency token designs, which have drawn legal scrutiny (U.S. Department of Justice, [n. d.]).
Advanced Anti-Auditing Techniques. Recent studies and case analyses have revealed various deceptive tactics involving fake ownership renunciation. Normally, when developers invoke renounceOwnership(), the contract should no longer have an administrator. However, malicious actors employ advanced techniques to retain control. Shiaeles and Li (2024) documented a real-world token case where misleading variable names were used within the contract. For example, the public variables owner and getOwner were set to the zero address, making ownership appear to have been renounced. Yet, a hidden control variable, such as the seemingly harmless isTokenReceiver, still stored the original developer’s address (Hall et al., 2024). As a result, although the contract appeared to have relinquished ownership, a fixed backdoor tied to a designated address remained. This type of fake renunciation is both highly covert and technically complex. Beyond ownership deception, many contracts use innocuous-sounding function names to disguise malicious logic. For instance, in the aforementioned case, the function name isTokenReceiver concealed administrative privileges. Similarly, other scam contracts may include functions with names like failsafe() or emergency(), which internally enable fund extraction or permission restoration, accessible only to the developer.
@@
MEV Bot Obfuscation Techniques. To protect MEV bots from being front-run by generalized mimicking scripts in the public mempool, practitioners and researchers have developed various obfuscation and privacy-preserving techniques. The most common method is using private relays (e.g., Flashbots) to submit bundles directly to block builders, bypassing the public mempool and preventing adversaries from copying transactions (Yang et al., 2024; Qin et al., 2022). Intent-based protocols like CoW Swap perform off-chain batch matching of user intents, publishing only the final settlement on-chain to eliminate front-running risks (Yüksel, [n. d.]). Gas camouflage techniques, such as locking transactions to specific tx.gasprice values or adding dummy computations, confuse adversarial repricing strategies (Yang et al., 2024). Multi-hop contract calls, often paired with flash loans and non-standard swap paths, increase attackers’ simulation overhead (Yüksel, [n. d.]; Daian et al., 2019). Bytecode-level obfuscations, like inserting JUMPI pseudo-branches or splitting constants via arithmetic, hinder static and dynamic analysis (Yüksel, [n. d.]). Recent work has also explored threshold encryption, delayed reveal schemes, and protocol-level MEV “tax” mechanisms to internalize ordering profits (Alnajjar et al., 2024; Daian et al., 2019). Despite these advancements, systematic research on obfuscation techniques for MEV bots at the smart contract level remains scarce.
10. Conclusion
@@
In this paper, we systematically investigate obfuscation techniques in Ethereum smart contract scams, providing comprehensive definitions, quantitative methods, and empirical analysis. We introduce a novel approach based on the detailed analysis of transfer instructions, identifying seven key quantifiable obfuscation features. Using a robust Z-score screening method, we analyze over 1.04 million Ethereum contracts, isolating approximately 3,000 (top 0.3%) highly suspicious contracts. Further quantitative financial impact analysis shows a pronounced disparity between obfuscated and non-obfuscated scam contracts: obfuscated contracts show an extreme financial extraction capability. We also demonstrate that obfuscation significantly undermines their effectiveness. Overall, our findings underscore the severe security risks posed by obfuscation and highlight the urgent need for advanced analytical and detection methodologies to address these evolving threats, enhancing blockchain security and fostering transparency.
Appendix A Extremely Centralized Contracts
A.1. Centralized Permission Control
(1) Overall Scam Logic Overview
@@ A.1. Centralized Permission Control
Such contracts often appear under the guise of ”lending protocols”, ”collateral management”, ”liquidity safeguarding”, etc., mimicking the interfaces and function names of well-known protocols like Compound or Uniswap. However, in reality, they are entirely controlled by a few roles—such as admin and liquidateAdmin (and sometimes more, e.g., manager or _super)—that manage all key operations.
@@
(2) Analysis of Code-Level Obfuscation Techniques
@@
  • Redundant Permission Checks: The contract repeatedly uses REQUIRE(msg.sender == admin, …) and REQUIRE(msg.sender == liquidateAdmin, …) in various locations. Since these checks are essentially equivalent, they further complicate code readability and give the false impression of a robust permission system.
@@
      21 For example, functions such as
      22 \_setMintGuardianPaused(), \_setBorrowGuardianPaused(), and \_setPauseGuardian() ostensibly provide multiple safeguard roles; however, a single admin instruction can control all permissions.
[…]
      24
@@ 21 For example, functions such as
      29function autoLiquidity(
      30address _account,
      31uint _liquidityAmount,
      32uint _liquidateIncome)
      33public returns (uint) {
      34REQUIRE(msg.sender == liquidateAdmin,
      35"SET_PRICE_ORACLE_OWNER_CHECK");
@@
      37// Actually calls
      38//autoLiquidityInternal(...)
@@
[…]
      45asset.ownerTransferToken(_owner,
      46_account, vars.aTokenBalance);
      47asset.ownerCompensation(_owner,
      48_account, vars.aTokenBorrow);’
@@
    • Redundant Functions: Functions such as _set-MintGuardianPaused(),_setBorrow-GuardianPau-sed(), _setTransferPaused(), autoLiquidityInternal(), an-d autoClearanceInternal() have nearly identical internal logic but are implemented in several different versions.
@@
(3) Core Features Identifiable from a Bytecode/Tool Perspective
@@
A.2. Unreasonable and Arbitrarily Adjustable High Fee / Tax Contracts
(1) Overall Scam Logic Overview
@@ A.2. Unreasonable and Arbitrarily Adjustable High Fee / Tax Contracts
(2) Code-level Obfuscation/Backdoor Technique Analysis
@@ 9 FeeOnTransferTokens(
(3) Core Features Extractable from Bytecode/Tool Detection
  • High Complexity in Transfer Logic: Within the _transfer() function, the frequent insertion of string operations and branch conditions results in elevated values for the features ”branch tree depth of address generation” and ”emit log density.”
@@
A.3. Contracts Without Genuine Fund Locking
(1) Overall Scam Logic Overview

### v2

@@ VII Impact on Existing Scam Detection Tools
Non-obfuscated 361 287 74 1.00 0.80 0.89
Obfuscated 92 11 81 1.00 0.12 0.21
Experimental Results. Table XVIII shows a large gap between non-obfuscated and obfuscated samples. For non-obfuscated contracts, SourceP achieves a precision of 1.00, a recall of 0.80, and an F1 of 0.89, which aligns with its intended capabilities. For obfuscated contracts, recall drops to 0.12 and F1 to 0.21 (precision remains high because the model predicts positives very conservatively), indicating that obfuscation sharply reduces detection effectiveness. High precision coupled with low recall implies conservative behavior: fewer false alarms but far more misses, which is undesirable for early warning.
The pronounced performance disparity between obfuscated and non-obfuscated samples can be attributed to how obfuscation disrupts SourceP’s static analysis pipeline. First, externalized transfer paths (T4) relocate the payable sink beyond the intra-contract scope that SourceP primarily analyzes, resulting in under-approximated money-flow graphs and the omission of critical Ponzi indicators. Second, opaque predicates and deep nesting (T2) complicate the control-flow graph, leading to path pruning or premature cut-offs during data and control-flow reconstruction, thereby suppressing essential features on which SourceP relies, such as cyclic payouts and balance-dependent branches. Third, multi-step address synthesis (T1), achieved through hashing, bit-masking, or nonces, disrupts constant propagation and complicates recipient attribution, thereby weakening heuristics that depend on recognizable payout targets. Fourth, dispatcher flattening and function cloning (T5) obscure function boundaries and hinder the reuse of summaries, which degrades feature aggregation at the function level. Fifth, log interference (T6) introduces spurious events and alters the order of informative logs, potentially misleading log-aware heuristics or subsequent sanity checks. Finally, junk code inflation (T7) increases bytecode size and the noise in the intermediate representation, thereby increasing the likelihood of timeouts or the application of conservative defaults. In contrast, non-obfuscated contracts reveal canonical money-flow structures that are consistent with SourceP’s feature design and training distribution, resulting in significantly enhanced effectiveness.
We qualitatively inspect representative misclassified obfuscated contracts and find that errors primarily occur in cases involving externalized transfer paths (T4) and path explosion (T2). Additionally, captions and logs often reveal partial analyses, such as truncated call-graph slices. The performance gap between obfuscated and non-obfuscated samples stems from reduced observability of Ponzi-specific signals under obfuscation, rather than from noisy labels or dataset differences. It is essential to note that our evaluation employs an ETH-only victim/flow definition, excluding intermediaries such as routers, centralized exchanges, and maximal extractable value (MEV). This means that the observed gap reflects limitations in our analysis rather than metric contamination.
@@
Answer to RQ4. Our evaluation results demonstrate that obfuscation has a significant detrimental impact on existing scam detection tools, causing a substantial drop in recall and F1. At a practical level, this observation underscores the importance of developing effective obfuscation analysis techniques to mitigate emerging security threats.
VIII Discussion
VIII-A Limitations and Threats to Validity
Our taxonomy (T1–T7) and features (F1–F7) focus on transfer-path obfuscation. Within this formal scope, the taxonomy aims for comprehensive coverage. Nonetheless, our work still has the following five limitations: 1) We do not cover ERC-20/721 token flows, bridge/mixer interactions, or metamorphic designs via CREATE2 or code replacement. These are outside our transfer-path scope and may require additional primitives. 2) The analysis is bytecode/IR–only (no dynamic execution or cross-transaction traces). Obfuscation that resolves purely at runtime (e.g., oracle-fed recipients, environment-gated dispatch) can reduce observability and yield false negatives. 3) When the payable path is fully delegated via DELEGATECALL/CALL to external contracts, F4 flags externalization but cannot recover the callee’s logic; downstream features (e.g., F1–F7) may thus remain unset. 4) Z-score and absolute guards rely on baseline distributions; different snapshots can shift thresholds slightly (we release seeds to recompute them). 5) Parts of the benchmarks (e.g., the SourceP comparison) use single-annotator manual labels; we exclude ambiguous cases and will release address lists, but residual noise may remain. 6) Our large-scale measurement focuses on the Ethereum mainnet due to its impact and data availability. We leave extending the study to other blockchains (e.g., BSC, Polygon, Arbitrum) for future work. Overall, we regard F1–F7 as a minimal, extensible core tailored to the transfer-path scope; extending to non-transfer flows or adding orthogonal runtime evidence is left to future work.
VIII-B Challenges and Future Work
Obfuscation techniques in smart contracts present significant challenges for auditing and regulatory practices, particularly in scam contracts, MEV bots, and highly centralized systems. These challenges include poor code readability, failure of conventional static detection tools, and delays in regulatory response. Several improvements are necessary to address these issues. Static analysis tools must be enhanced to track deeper control and data flows, while de-obfuscation preprocessing can simplify bytecode for more efficient audits. Dynamic analysis, such as runtime tracing or fuzzing, can bypass superficial obfuscation and validate fund flows during contract execution. Additionally, techniques from traditional software security, such as CFG flattening and semantic normalization, can be applied to EVM bytecode to identify critical logic, such as transfers and permission checks, that obfuscation attempts to conceal. Eventually, establishing collaborative platforms (e.g., a ”contract blacklist” or a ”high-risk obfuscation” repository) would enable researchers, auditors, and the public to tag suspicious contracts, thereby improving transparency and enhancing collective oversight of the DeFi ecosystem.
VIII-C Obfuscation signals vs. malicious intent
Our taxonomy (T1–T7) and features (F1–F7) capture how observable the transfer logic is, not whether that logic is benign or harmful. The same primitives often appear in legitimate designs. For instance, T4 can be used in upgradeable proxies and routers, while T5 in flattened dispatchers for gas efficiency. Hence, we treat F1–F7 as signals that warrant scrutiny, not verdicts. When we discuss scams, the goal is to illustrate correlation, not to claim causation. Disambiguating intent typically requires orthogonal evidence (economic semantics, longitudinal flows, victim-side signals), which is outside the scope of our obfuscation quantification.
IX Related Work
Classical Obfuscation Techniques. A large software engineering (SE) literature has studied source/IR-level obfuscation for native/managed code. Collberg et al. offer a widely used taxonomy (layout, control-, data-, and preventive transformations) with potency/resilience/stealth metrics [11]. Representative techniques include opaque predicates and bogus control-flow (control-flow obfuscation), anti-disassembly/anti-decompilation (preventive), instruction substitution and data encoding, and control-flow flattening; industrial-strength implementations such as Obfuscator-LLVM operationalize several of these passes [35, 31, 29]. Surveys and monographs synthesize two decades of progress and limitations [53, 10].
@@ IX Related Work
Smart Contract Security Analysis. Some research employs static analysis to enhance the security and efficiency of smart contracts. USCHUNT [4] explores the balance between adaptability and security in upgradeable contracts. Madmax [22] targets vulnerabilities to prevent execution failures, while Slither [56] and Smartcheck [62] automatically detect flaws in Solidity contracts. Symbolic execution is also used to improve security; Mythril [13] analyzes EVM bytecode, EthBMC [20] combines symbolic execution with concrete validation, and Reguard [37] and Manticore [41] identify reentrancy and other bugs. Smartian [8] integrates fuzzing with static and dynamic analysis, while Confuzzius [19] leverages data dependency insights for fuzzing. ContractFuzzer [28] and Sfuzz [40] apply fuzzing to uncover security issues. Research highlights various formal verification methods to enhance the security of smart contracts. Sailfish [5] improves state inconsistency detection, while VetSC.[18] extends DApp verification. Zeus[30] and Verx [43] focus on contract safety and condition verification. Smartpulse [59] analyzes time-based properties, Securify [63] identifies security breaches, and Verismart [57] ensures contract safety.
Scam and Ponzi detection. Prior work has studied Ponzi detection using code-level features and runtime behavior graphs, including SADPonzi [7], SourceP [38], and PonziGuard [34]. These approaches are effective on non-obfuscated or lightly obfuscated contracts but are sensitive to transfer-path obfuscation and externalized payout logic. Very recent work such as MFDPonzi [6] explores static features from opcode sequences at larger scale; on our labeled set, MFDPonzi shows comparable precision/recall trends to SourceP, so we use SourceP as a representative baseline in the main evaluation.
Advanced Anti-Auditing Techniques. Recent analyses expose “fake” ownership renunciation: after invoking renounceOwnership(), some contracts zero public fields (e.g., owner, getOwner) yet retain control via concealed state. Shiaeles and Li (2024) document a live case where a benign-looking variable (e.g., isTokenReceiver) stores the deployer’s address, leaving an address-bound backdoor [25]. More broadly, innocuous names such as isTokenReceiver, failsafe(), or emergency() can mask administrator-only operations (e.g., privilege restoration or fund extraction).
@@
MEV Bot Obfuscation Techniques. To protect MEV bots from being front-run by generalized mimicking scripts in the public mempool, practitioners and researchers have developed various obfuscation and privacy-preserving techniques. The most common method is using private relays (e.g., Flashbots) to submit bundles directly to block builders, bypassing the public mempool and preventing adversaries from copying transactions [70, 44]. Intent-based protocols like CoW Swap perform off-chain batch matching of user intents, publishing only the final settlement on-chain to eliminate front-running risks [70]. Gas camouflage techniques, such as locking transactions to specific tx.gasprice values or adding dummy computations, confuse adversarial repricing strategies [70]. Multi-hop contract calls, often paired with flash loans and non-standard swap paths, increase attackers’ simulation overhead [70, 17]. Bytecode-level obfuscations, like inserting JUMPI pseudo-branches or splitting constants via arithmetic, hinder static and dynamic analysis [70]. Recent work has also explored threshold encryption, delayed reveal schemes, and protocol-level MEV “tax” mechanisms to internalize ordering profits [3, 17]. Despite these advancements, systematic research on obfuscation techniques for MEV bots at the smart contract level remains scarce.
X Conclusion
@@
In this paper, we systematically investigate obfuscation techniques in Ethereum smart contracts, providing comprehensive definitions, quantitative methods, and empirical analysis. We introduce seven key quantifiable obfuscation features based on the detailed analysis of transfer instructions. Using a robust Z-score representation model, we analyze the prevalence of obfuscation techniques employed in over 1.04 million Ethereum contracts and conduct an in-depth analysis of 3,000 highly suspicious contracts, revealing four types of malicious contracts. Our further analysis shows that obfuscated scam contracts have a higher financial extraction capability than non-obfuscated scam contracts. We also demonstrate that obfuscation significantly undermines the effectiveness of existing detection tools. At ecosystem scale, obfuscation is routine but risk concentrates in the extreme tail; in practice, auditing should prioritize high-score segments and signals such as F3/F2/F4 for tiered review. Overall, our findings underscore the security risks posed by obfuscation and highlight the urgent need for advanced analytical and detection methodologies to address this evolving threat, enhancing blockchain security and fostering transparency.
Supplementary Overview
This supplementary material provides code-level evidence for the three extreme centralization patterns discussed in the main paper (Case IV: Extreme Centralization). We organize the evidence into S1–S3: centralized permission control, arbitrarily adjustable high fee/tax, and lack of genuine fund locking. Each subsection follows a unified structure (summary, code evidence, and tool-visible signals) to improve scanability.
Appendix A Supplementary Evidence for Extreme Centralization (S1–S3)
A-A S1: Centralized Permission Control
Summary.
@@ A-A S1: Centralized Permission Control
Such contracts often appear under the guise of ”lending protocols”, ”collateral management”, ”liquidity safeguarding”, etc., mimicking interfaces and function names of well-known protocols (e.g., Compound, Uniswap). However, in reality, they are entirely controlled by a few roles—such as admin and liquidateAdmin (and sometimes more, e.g., manager or _super)—that manage all key operations.
@@
Code evidence.
@@
  • Redundant Permission Checks and Guard Inflation. The contract repeats equivalent REQUIRE(...) guards for admin and liquidateAdmin, inflating complexity without adding real protection.
@@ 25 $\begin{applist}$
      27 $\begin{lstlisting}$[language=Solidity]
      28function autoLiquidity(
      29address _account,
      30uint _liquidityAmount,
      31uint _liquidateIncome)
      32public returns (uint) {
      33REQUIRE(msg.sender == liquidateAdmin,
      34"SET_PRICE_ORACLE_OWNER_CHECK");
@@ 27 $\begin{lstlisting}$[language=Solidity]
      36// Actually calls
      37//autoLiquidityInternal(...)
@@
      43 $\begin{lstlisting}$[language=Solidity]
      44asset.ownerTransferToken(_owner,
      45_account, vars.aTokenBalance);
      46asset.ownerCompensation(_owner,
      47_account, vars.aTokenBorrow);’
@@ 43 $\begin{lstlisting}$[language=Solidity]
    • $\circ$
      Redundant Functions: Functions such as _setMintGuardianPaused() , _setBorrowGuardianPaused() , _setTransferPaused() , autoLiquidityInternal() , and autoClearanceInternal() have nearly identical internal logic but are implemented in several different versions.
    • $\circ$
@@
Tool-visible signals.
@@
A-B S2: Unreasonable and Arbitrarily Adjustable High Fee / Tax Contracts
Summary.
@@ A-B S2: Unreasonable and Arbitrarily Adjustable High Fee / Tax Contracts
Code evidence.
@@ 9 FeeOnTransferTokens(
Tool-visible signals.
  • High Complexity in Transfer Logic. Within the _transfer() function, frequent string operations and branch conditions elevate the branch-depth and log-density signals.
@@
A-C S3: Contracts Without Genuine Fund Locking
Summary.